OpenStudy (anonymous):

Solve for x: (x^4 - 1) / (x^3) = 0

OpenStudy (anonymous):

x=-1 x=1

OpenStudy (dumbcow):

denominator is irrelevant, since zero divided by anything is still 0 x^4 - 1 = 0 x^4 = 1 x = +-1

OpenStudy (anonymous):

Thanks..can u help me with this: i was trying to solve it using the quadratic formula. im not sure if its rite thou

OpenStudy (dumbcow):

sure

OpenStudy (anonymous):

solve for x: 2x^2 +5x = 8

OpenStudy (dumbcow):

let a = 2, b=5, c = -8 \[x = \frac{-5 \pm \sqrt{(-5)^{2}-4(2)(-8)}}{2(2)}\]

OpenStudy (dumbcow):

\[x = \frac{-5 \pm \sqrt{89}}{4}\]

OpenStudy (anonymous):

Isnt teh quadratic formula x = -b +/ sqrt b^2 not (-b)^2

OpenStudy (anonymous):

so shudnt it be jsut 5^2 not (-5)^2

OpenStudy (anonymous):

never mind...i get it cuz it makes no dffnce

OpenStudy (anonymous):

teh outcoem is teh same

OpenStudy (anonymous):

@arizona cardianl can u help me with a problem..its long and confusing?????

OpenStudy (dumbcow):

yesh sure, and you are right i messed up but since its squared it makes no difference good catch though :)

OpenStudy (anonymous):

Solve for x: (x+1)^2 (x-2) + (x+1)(x-2)^2 = 0

OpenStudy (dumbcow):

factor out what they have in common each term has a (x+1)(x-2) \[(x+1)(x-2)((x+1)+(x-2)) = 0\] \[(x+1)(x-2)(2x-1) = 0\]

OpenStudy (anonymous):

oh ok so then teh solutions are x = -1, x = 2, x = 1/2

OpenStudy (dumbcow):

yep make sense now

OpenStudy (anonymous):

yes..i thought it was going to be harder because there were exponents.

OpenStudy (anonymous):

i forgot absolute values: so the q askes\[\left| x-3 \right|<7 \] when we solve it do we have to set it to (x-3) is less tahn 7 and teh otehr one to (x-3) greater tahn -7?

OpenStudy (dumbcow):

yes that is correct \[-7 < x-3 < 7\]